Alternative Inventory Methods Park Company's perpetual inventory records indicate the following transactions in the month of...

  1. Alternative Inventory Methods

    Park Company's perpetual inventory records indicate the following transactions in the month of June:

    Units Cost/Unit
    Inventory, June 1 200 $3.20
    Purchases:
          June 3 200 3.50
          June 17 250 3.60
          June 24 300 3.65
    Sales:
          June 6 300
          June 21 200
          June 27 150

    Required:

    1. Compute the cost of goods sold for June and the inventory at the end of June using each of the following cost flow assumptions: If required, round your answers to the nearest dollar.
    1. FIFO
      Cost of Goods Sold $
      Ending Inventory $
    2. LIFO (Round your intermediate calculations and final answers to the nearest cent.)
      Cost of Goods Sold $
      Ending Inventory $
    3. Average cost (In your computations, round new per unit costs to the nearest cent. Round your intermediate computations and final answers to the nearest dollar.)
      Cost of Goods Sold $
      Ending Inventory $
Answer #1
  1. FIFO
    Cost of Goods Sold $1095
    Ending Inventory $2240
  2. LIFO (Round your intermediate calculations and final answers to the nearest cent.)
    Cost of Goods Sold $1047.50
    Ending Inventory $2287.50
  3. Average cost (In your computations, round new per unit costs to the nearest cent. Round your intermediate computations and final answers to the nearest dollar.)
    Cost of Goods Sold $1083
    Ending Inventory $2252

Explanations :-

FIFO :-

Total units available for sale = Beginning inventory + Purchase

= 200 + 200 + 250 + 300 = 950 units

Cost of Goods available for sale = (200 * $3.20) + (200 * $3.5) + (250 * $3.6) + (300 * 3.65) = $3335

Ending Inventory consists 300 units purchased as on June 24

= 300 units * $3.65 = $1095

Cost of goods sold = Cost of Goods available for sale - Ending Inventory

= $3335 - $1095 = $2240

LIFO :-

Ending Inventory consists 100 units from Beginning Inventory, 50 units from June 17 & 150 units from June 24

= (100 * $3.20) + (50 * $3.60) + (150 * $3.65) = $1047.50

Cost of goods sold = Cost of Goods available for sale - Ending Inventory

= $3335 - $1047.50 = $2287.50

Average :-

(200 * $3.20) + (200 * $3.50)/400 = $3.35

Units sold = 300 units

Remaining = 100 units @ $3.35

Purchase June 17 = (250 * $3.60) + (100 * $3.35)/350 = $3.53

Sale June 21 = 200 units

Remaining units = 150 units @ $3.53

Purchase June 24 = (300 * $3.65) + (150 * $3.53)/450 = $3.61

Sale June 27 = 150 units

Ending Inventory = 300 units * $3.61 = $1083

Cost of goods sold = $3335 - $1083 = $2252